New microelectronics testing equipment was purchased 2 years ago by Mytesmall Industries at a cost of $600,000. At that time, it was expected to be used for 5 years and then traded or sold for its salvage value of $75,000. Expanded business in newly developed international markets is forcing the decision to trade now for a new unit at a cost of $800,000. The current equipment could be retained, if necessary, for another 2 years, at which time it would have a $5000 estimated market value. The current unit is appraised at $350,000 on the international market, and if it is used for another 2 years, it will have M&O costs (exclusive of operator costs) of $125,000 per year. Determine the values of P, n, S, and AOC for this defender if a replacement analysis were performed today. Make without using excel and detail.  
Answer #1

If the current unit is retained for another two years:

Annual M&O cost=$125,000

Salvage value after 2 years=$5,000

Assuming discount rate of 10%=0.1

Present value(PV) of annual M& O cost:

Present value(PV)of cash flow=(Cash flow)/((1+i)^N)

N=year of cash flow, i=discount rate

PV of M& O cost=(125000/1.1)+(125000/(1.1^2))=$113,636.36+ $ 103,305.79 = $ 216,942.15

PV of Salvage value=5000/(1.1^2)= $ 4,132.23

Net Present Worth of cost for the defender=P=$216,942.15- $ 4,132.23 = $ 212,809.92

n=2 years

Capital Recovery Factor (A/P,I,n)=CRF=(i((i+1)^n))/(((1+i)^n)-1)

CRF=(0.1*(1.1^2))/((1.1^2)-1)= 0.576190476

AOC=Annual Operating Cost=P*CRF=0.576190476*($ 212,809.92)= $ 122,619.05

The challenger (New unit) has a Present Value of cost=$800,000-$350,000=$$450,000

Assuming economic life of 5 years :

N=5 years

CRF =(0.1*(1.1^5))/((1.1^5)-1)= 0.263797481

AOC=450000*0.263797481= $ 118,708.87